# Structure-Preserving Transformers for Sequences of SPD Matrices

### 14 Dec 2023 v4

@@ Abstract
Index Terms—  Transformers, SPD Matrices, Structure-Preserving, Electroencephalography, Sleep Staging
1 Introduction
@@
ptQptptptLinearptptptLinearptKptMatmulptSoftmaxptLinear CombinationptMatmulptVpt$(N\times L\times d)$$(N\times L\times d)$$(N\times L\times L)$$h$$\begin{array}[]{c}([N\times h]\times L\\ \newline \times[d/h])\end{array}$ $([N\times h]\times L\times L)$$(N\times L\times d)$
Fig. 1: The SP-MHA architecture. In parentheses are tensor dimensions at every step, with $N$ the batch size.
Fig. 2: SPDTransNet global architecture, with $t=3$ feature tokens per epoch.
@@
2 SPD Structure-Preserving Attention
@@
2.1 Structure-Preserving Multihead Attention (SP-MHA)
@@
with $\mathcal{L}_{Q}(\cdot)$ and $\mathcal{L}_{K}(\cdot)$ banks of $h$ linear maps from $\mathbb{R}^{d(m)}$ to $\mathbb{R}^{\frac{d(m)}{h}}$, $sm(\cdot)$ the softmax function, and $C(\cdot)$ the weighted linear combination of the $h$ post-softmax attention maps. Although said attention maps are identical to their L-MHA counterparts, the only operation applied to V is the final matrix multiplication, i.e. linear combinations of V’s tokens weighted by the combined attention map, which do not threaten our tokens’ vector space geometry.
2.2 Triangular linear maps
@@
Table 1: Results obtained from both our model and the re-trained literature. Best results are in bold.
3 Application to EEG Sleep Staging
@@
As seen in a recent survey by Phan et al. [17], state-of-the-art automatic sleep staging models typically use two-step architectures - given a sequence of epochs, epoch-wise features are extracted before being compared at the sequence-wise level, utilizing this contextual information to improve classification. Since epochs often contain markers indicative of multiple stages, two-step architectures tend to subdivide them further, extracting features from subwindows using convolutional NNs [12] and/or recurrent NNs [18, 19] - the latter utilizing RNNs for both steps. Multiple authors have adapted this context-inclusive approach to Transformer-based architectures [20, 21, 15], with auto-attention mechanisms at both the intra- and inter-epoch levels, taking advantage of the high performance they offer when applied to sequence-based data.
3.1 Improving stage-wise classification
@@
According to the aforementioned survey [17], current sleep staging models have attained a sufficient performance level to replace manual staging in some contexts. However, we have found that class-wise performance was often lacking, particularly with regards to the N1 sleep stage [16], universally difficult to classify (Section 4). Most EEG datasets are heavily imbalanced, with the N1 stage often underrepresented (Section 3.3) - models optimized for high overall accuracy may thus sacrifice N1 classification if it improves global performance. To account for this, recent approaches [14, 15] elected to primarily evaluate their performance through the macro-averaged F1 (MF1) score, a class-wise balanced metric widely used in the literature. They also rebalance their training sets through oversampling, so that all stages within have the same number of classification targets. While the survey states that a sequence-to-sequence classification scheme (classifying each epoch in the input sequence) might lead to better performance, having multilabel inputs is nonsensical for this rebalancing - hence their use of a sequence-to-epoch scheme (classifying one epoch per sequence).
@@
Seraphim et al. [15] hypothesized that an analysis through functional connectivity - the activation correlations between different brain regions [22] - enhances stage-wise performance. Such an analysis was first done by Jia et al. [13], using epoch-wise graph learning to estimate said connectivity and sequence-wise spatio-temporal graph NNs to compare them. By contrast, Seraphim et al. estimate it through covariance matrices. Their two-step model uses standard Transformer encoders at each step, reminiscent of [21]. Each input epoch is described as a multichannel timeseries of SPD matrices, which are then tokenized bijectively. However, their approach does not guarantee the preservation of their data’s SPD structure, as they operate a channel-wise concatenation of their tokens, in addition to the concatenations found within their encoders (Section 2.1). Hence, we propose a Transformer-based model capable of analyzing EEG-derived functional connectivity through SPD matrices without sacrificing the SPD structure of our data throughout the analysis.
3.2 The SPDTransNet model
@@
Each EEG signal is decomposed into $C$ channels, divided into epochs, and further subdivided into $S$ subwindows per epoch. After preprocessing (Section 3.3), each epoch is described by $S\times C$ matrices in $SPD(n)$. Each matrix is mapped onto $Sym(n)$ logarithmically (Section 1), tokenized (Section 2), and linearly mapped onto $Sym(m)$ (with $m>n$, as we have found that larger tokens improve performance). The $S\times C$ grid of tokens is then arranged into a 1D sequence, with the $S$ tokens in the channel 1 followed by the $S$ tokens in channel 2, etc.
@@
We ensure structure preservation by using the SP-MHA bloc in all Transformer encoders, and choosing all linear maps within said encoders’ Feed-Forward (FF) components [5] and the aforementioned FC blocs to be triangular (Section 2.2). The ReLU and dropout layers in the FF and FC blocs do not cause issue, as setting a values within a token to 0 won’t remove the corresponding matrix from $Sym(m)$. Same for the positional encodings, average poolings and in-encoder layer normalizations, which all qualify as linear combinations.
@@ 4 Experiments & Results
To maximize class-wise performance, we operate a hyperparameter research per configuration, followed by a 31-fold cros-validation. As do [14, 15] (Section 3.1), we rebalance all training sets and maximize the MF1 score. To explore the importance of the context length $\ell$ (Section 3.2) within our model, we ran hyperparameter researches with $\ell$ = 6, 10 or 14 (i.e. $L$ = 13, 21 or 29), with hyperparameter research configuration unchanged between them.
@@
Our hyperparameter researches use the Optuna tool [24], with 5 simultaneous runs and 50 total runs per configuration. Hyperparameters include\@footnotemark the token size $d(m)$, set by the first linear map (Section 3.2) and chosen in {351, 378} (i.e. $m$in {26, 27}) [^3] Since $\frac{d(m)}{h}$ must be an integer, potential values for those are limited.; the $h$ parameter of each Transformer encoder, in {3, 9}\@footnotemark; and the number of epoch feature tokens $t$ (Section 3.2), chosen among {1, 3, 5, 7, 10} - with in particular $t=1$ akin to describing each epoch with a single token, and $t=7$ corresponding to one token being preserved per channel. We train all folds on the hyperparameters giving the best validation MF1, as well as those with the best F1 score for the N1 stage. Out of those two sets, the results from the set yielding the best average test MF1 is presented in lines 6 to 8 of Table 1, with the corresponding hyperparameter set, $d(m)$ and $t$ in the final three columns.
We compare ourselves to five models: DeepSleepNet [11], often used as a benchmark, with a pre-trained epoch-wise global feature map submodel followed by a sequence-to-sequence RNN; IITNet [12], the source of our 31 folds, extracting multiple features per epoch through CNNs and comparing them through sequence-wise RNNs; GraphSleepNet [13], expliciting epoch-wise functional connectivity through graph learning; Dequidt et al. [14], utilizing a single-step pretrained visual CNN, who both maximize MF1 performance and rebalance training sets; and Seraphim et al. [15], with a similar approach to ours lacking in structural preservation (Section 3.1). These models were re-trained using our methodology - except for oversampling in DeepSleepNet’s sequence-to-sequence submodel - though we use only their published hyperparameters. Finally, as test sets vary between models due to sequence-based recording-wise border effects, we trim test set borders to enforce uniformity. All these changes cause the results we obtained to differ somewhat from those initially published. Our results, averaged over all folds, are displayed in lines 1 to 5 of Table 1.
@@
As shown in Table 1, we obtain the best MF1 and N1 F1 scores for $L=21$, whereas the best macro-averaged accuracy is obtained for $L=13$. For all values of $L$, we outperform the state-of-the-art on the considered metrics (except for the MF1 score for $L=29$). Moreover, all three configurations have around a two-point lead in both macro accuracy and N1 F1 score. While our model favors the smaller token size of $d(m)=351$ for all values of $L$, it seems that having a large number of tokens to describe each epoch (at least $t=5$) is necessary for best performance. Overall, $L=21$ seems to be a good compromise to capture enough contextual information without burdening our model with irrelevant data.
5 Conclusion

### 06 Mar 2024 v5

# ENSICAEN, CNRS, GREYC

@@ Abstract
Index Terms:
Transformers, SPD Matrices, Structure-Preserving, Electroencephalography, Sleep Staging
I Introduction
@@
Figure 1: The SP-MHA architecture. In parentheses are tensor dimensions at every step, with $N$ the batch size.
Figure 2: SPDTransNet global architecture, with $t=3$ feature tokens per epoch.
@@
II SPD Structure-Preserving Attention
@@
II-A Structure-Preserving Multihead Attention (SP-MHA)
@@
with $\mathcal{L}_{Q}(\cdot)$ and $\mathcal{L}_{K}(\cdot)$ banks of $h$ linear maps from $\mathbb{R}^{d(m)}$ to $\mathbb{R}^{\frac{d(m)}{h}}$, $sm(\cdot)$ the softmax function, and $C(\cdot)$ the weighted linear combination of the $h$ post-softmax attention maps. Here, the computation of attention maps (small-dashed black rectangle in Figure 1) remains identical to L-MHA. However, their application to V (large-dashed red rectangle in the figure) only requires a matrix multiplication, i.e. linear combinations of V’s tokens weighted by the combined attention map. As such, the SP-MHA bloc does not threaten our tokens’ vector space geometry.
II-B Triangular linear maps
@@
TABLE I: Results obtained from both our model and the re-trained literature. Best results are in bold.
III Application to EEG Sleep Staging
@@
As seen in a recent survey by Phan et al. [17], state-of-the-art automatic sleep staging models typically use two-step architectures - given a sequence of epochs, epoch-wise features are extracted before being compared at the sequence-wise level, utilizing this contextual information to improve classification. Since epochs often contain markers indicative of multiple stages, two-step architectures tend to subdivide them further, extracting features from subwindows using convolutional NNs [11, 12] and/or recurrent NNs [18, 19, 20] - the latter utilizing RNNs for both steps. Multiple authors have adapted this context-inclusive approach to Transformer-based architectures [21, 22, 15], with auto-attention mechanisms at both the intra- and inter-epoch levels, taking advantage of the high performance they offer when applied to sequence-based data.
III-A Improving stage-wise classification
@@
According to the aforementioned survey [17], current sleep staging models have attained a sufficient performance level to replace manual staging in some contexts. However, we have found that class-wise performance was often lacking, particularly with regards to the N1 sleep stage [16], universally difficult to classify (Section IV). Most EEG datasets are heavily imbalanced, with the N1 stage often underrepresented (Section III-C) - models optimized for high overall accuracy may thus sacrifice N1 classification if it improves global performance. To account for this, recent approaches [14, 15] elected to primarily evaluate their performance through the macro-averaged F1 (MF1) score, a class-wise balanced metric widely used in the literature. They also rebalance their training sets through oversampling, so that all stages within have the same number of classification targets. While the survey states that a sequence-to-sequence classification scheme (classifying each epoch in the input sequence) might lead to better performance, having multilabel inputs is nonsensical for this rebalancing - hence their use of a sequence-to-epoch scheme (classifying one epoch per sequence).
@@
Seraphim et al. [15] hypothesized that an analysis through functional connectivity - the activation correlations between different brain regions [23] - enhances stage-wise performance. Such an analysis was first done by Jia et al. [13], using epoch-wise graph learning to estimate said connectivity and sequence-wise spatio-temporal graph NNs to compare them. By contrast, Seraphim et al. estimate it through covariance matrices. Their two-step model uses standard Transformer encoders at each step, reminiscent of [22]. Each input epoch is described as a multichannel timeseries of SPD matrices, which are then tokenized bijectively. However, their approach does not guarantee the preservation of their data’s SPD structure, as they operate a channel-wise concatenation of their tokens, in addition to the concatenations found within their encoders (Section II-A). Hence, we propose a Transformer-based model capable of analyzing EEG-derived functional connectivity through SPD matrices without sacrificing the SPD structure of our data throughout the analysis.
III-B The SPDTransNet model
@@
Each EEG signal is decomposed into $C$ channels, divided into epochs, and further subdivided into $S$ subwindows per epoch. After preprocessing (Section III-C), each epoch is described by $S\times C$ matrices in $SPD(n)$. Each matrix is mapped onto $Sym(n)$ logarithmically (Section I), tokenized (Section II), and linearly mapped onto $Sym(m)$ (with $m>n$, as we have found that larger tokens improve performance). The $S\times C$ grid of tokens is then arranged into a 1D sequence, with the $S$ tokens in the channel 1 followed by the $S$ tokens in channel 2, etc.
@@
We ensure structure preservation by using the SP-MHA bloc in all Transformer encoders, and choosing all linear maps within said encoders’ Feed-Forward (FF) components [5] and the aforementioned FC blocs to be triangular (Section II-B). The ReLU and dropout layers in the FF and FC blocs do not cause issue, as setting a values within a token to 0 won’t remove the corresponding matrix from $Sym(m)$. Same for the positional encodings, average poolings and in-encoder layer normalizations, which all qualify as linear combinations.
@@ IV Experiments & Results
To maximize class-wise performance, we operate a hyperparameter research per configuration, followed by a 31-fold cros-validation. As do [14, 15] (Section III-A), we rebalance all training sets and maximize the MF1 score. To explore the importance of the context length $\ell$ (Section III-B) within our model, we ran hyperparameter researches with $\ell$ = 6, 10 or 14 (i.e. $L$ = 13, 21 or 29), with hyperparameter research configuration unchanged between them.
@@
Our hyperparameter researches use the Optuna tool [25], with 5 simultaneous runs and 50 total runs per configuration. Hyperparameters include\@footnotemark the token size $d(m)$, set by the first linear map (Section III-B) and chosen in {351, 378} (i.e. $m$in {26, 27}) [^3] Since $\frac{d(m)}{h}$ must be an integer, potential values for those are limited.; the $h$ parameter of each Transformer encoder, in {3, 9}\@footnotemark; and the number of epoch feature tokens $t$ (Section III-B), chosen among {1, 3, 5, 7, 10} - with in particular $t=1$ akin to describing each epoch with a single token, and $t=7$ corresponding to one token being preserved per channel. We train all folds on the hyperparameters giving the best validation MF1, as well as those with the best F1 score for the N1 stage. Out of those two sets, the results from the set yielding the best average test MF1 is presented in lines 1 to 3 of Table I, with the corresponding hyperparameter set, $d(m)$ and $t$ in the final three columns.
@@
As shown in Table I, we obtain the best MF1 and N1 F1 scores for $L=21$, whereas the best macro-averaged accuracy is obtained for $L=13$. For all values of $L$, we outperform the state-of-the-art on the considered metrics (except for the MF1 score for $L=29$). Moreover, all three configurations have around a two-point lead in both macro accuracy and N1 F1 score. While our model favors the smaller token size of $d(m)=351$ for all values of $L$, it seems that having a large number of tokens to describe each epoch (at least $t=5$) is necessary for best performance. Overall, $L=21$ seems to be a good compromise to capture enough contextual information without burdening our model with irrelevant data.
We also investigate the impact of our strict structural preservation by replacing the SP-MHA bloc of SPDTransNet model with the classic L-MHA (Section 2), all other things being equal (with $L=21$). Results for this configuration are displayed in line 4 of the table.
We compare ourselves to five models: DeepSleepNet [11], often used as a benchmark, with a pre-trained epoch-wise global feature map submodel followed by a sequence-to-sequence RNN; IITNet [12], the source of our 31 folds, extracting multiple features per epoch through CNNs and comparing them through sequence-wise RNNs; GraphSleepNet [13], expliciting epoch-wise functional connectivity through graph learning; Dequidt et al. [14], utilizing a single-step pretrained visual CNN, who both maximize MF1 performance and rebalance training sets; and Seraphim et al. [15], with a similar approach to ours, though utilizing an alternative whitening (Section III-C) and - similarly to line 4 of the table - lacking in structural preservation (Section III-A). These models were re-trained using our methodology - except for oversampling in DeepSleepNet’s sequence-to-sequence submodel - though we use only their published hyperparameters. Finally, as test sets vary between models due to sequence-based recording-wise border effects, we trim test set borders to enforce uniformity. All these changes cause the results we obtained to differ somewhat from those initially published. These results, averaged over all folds, are displayed in lines 5 to 9 of Table I. As we can see, SPDTransNet outperforms all tested State-of-the-Art models, though our lead on Dequidt et al. is minor.
Furthermore, comparing our best results (line 2) to those of lines 4 and 9 indicate that the structural preservation of our SP-MHA improves our model’s performance, with or without the influence of our new whitening (Section III-C).
V Conclusion